OpenStudy (anonymous):

How would I write y = 2x - 14 + 3? The answer is supposed to be y = 2x - 11

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

How would I write this

OpenStudy (anonymous):

y = 2x - 14 + 3

OpenStudy (anonymous):

y = 2x - 14 + 3 is y = 2x - (-11)

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@thomaster

hero (hero):

-14 + 3 = -11

OpenStudy (anonymous):

In slope - intercept form

OpenStudy (anonymous):

And the 11 is not supposed to be negative

OpenStudy (anonymous):

So I have no clue on how to solve this

hero (hero):

You don't write y = 2x - 14 + 3 you write y = 2x - 11 @jdoe001 just didn't want to give it away.

OpenStudy (anonymous):

Oh :(

OpenStudy (anonymous):

THanks anyway :D

OpenStudy (anonymous):

Medal for you :3